OpenStudy (anonymous):

In Pensacola in June, high tide was at noon. The water level at high tide was 12 feet and 2 feet at low tide. Assuming the next high tide is exactly 12 hours later and that the height of the water can be modeled by a cosine curve, find an equation for water level in June for Pensacola as a function of time (t).

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

can any one please help me solve this step by step?

OpenStudy (jpes2193):

cant you create a chart to find d=rt?

OpenStudy (anonymous):

what ? im sorry i dont understand :/

OpenStudy (jpes2193):

umm hold on. let me make sure Im giving you the right information

OpenStudy (anonymous):

ok

OpenStudy (jpes2193):

Ive not worked with cos functions. im sorry Im not going to be any help here.

OpenStudy (jpes2193):

@amistre64 might be able to help

OpenStudy (anonymous):

ok alright thanks

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

can one of you guys walk me through this?

OpenStudy (amistre64):

oh ive tried helping before :) we need to determine amplitude, and period to start things off with

OpenStudy (amistre64):

how do we define amplitude?

OpenStudy (anonymous):

amplitude is the highest point on the y axis which is the distance right? it has to be a positive number?

OpenStudy (amistre64):

amplitude is half the distance between the lowest and highest values for the Y axis. (high - low)/2 = amplitude from the stated problem what is the amplitude?

OpenStudy (amistre64):

these are the parts we want to determine. midline period vv vv y = M + A cos( P t) ^^ amp

OpenStudy (amistre64):

midline is the average of low and high M = (low+high)/2 amplitude is half the height, and M is in the center, so adding M+A gives us the high point A = high - M period is a measure of how long it takes to repeat itself, since cos(u) has a normal period of 2pi, and we know that it takes 12 hours to go to low tide, and itll presumably take another 12 hours to get back to high tide (general assumptions); then we t=24 .. u = Pt; when t=24, u=2pi P(24) = 2pi must be satisfied.

OpenStudy (anonymous):

i know that the amplitude is 5

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

good, and the midline?

OpenStudy (amistre64):

A = high - M we know A and high, find M or another way, average high and low

OpenStudy (anonymous):

what is the midline again?

OpenStudy (amistre64):

im not going to make a long drawn out posting of repeating myself over and over when what ive already typed is sufficient. scroll up and read. if its unclear, then ask questions about what ive already stated.

OpenStudy (anonymous):

ohhhh is it 7 ?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

now your going to make me ... 12 + 2 14 ... yeah 7s fine :)

OpenStudy (amistre64):

so far we have: y = 7 + 5cos(u) just have to determine u = Pt for the period ...

OpenStudy (anonymous):

so, 24=2pi do i simplify this? or factorize it ?

OpenStudy (amistre64):

24 P = 2pi, its a basic algebraic process to find P ...

OpenStudy (anonymous):

do i divide by 24 at both sides to isolate the variable?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

so P= 2pi/24

OpenStudy (amistre64):

or simply P = pi/12 now put the equation together y = M + A cos(Pt)

OpenStudy (anonymous):

do i simplify this?

OpenStudy (anonymous):

but pi over twelve isnt any of my options

OpenStudy (anonymous):

@amistre64 isnt there a way ti simplify it more?

OpenStudy (anonymous):

it*

OpenStudy (amistre64):

i dont know what your options are

OpenStudy (anonymous):

here they are

OpenStudy (anonymous):

t it be c?

OpenStudy (amistre64):

it takes 12 hours to go from high to low, so it would take 24 hours to go from high to high; P 24 = 2pi when P = pi/12

OpenStudy (amistre64):

Assuming the next high tide is exactly 12 hours later i misread the info ..... pi/6 is fine

OpenStudy (amistre64):

C is the most likely option since it fits the Midline and Amp as determined .... and pi/6 is good is correct when we do the actual problem instead of making things up lol

OpenStudy (anonymous):

XD thanks a lot for your help :-)

OpenStudy (amistre64):

your welcome :)